OpenStudy (anonymous):

Which value of a would make the following statement true -2(x + a) = -2x - 4? A. -4 B. -2 C. 2 D. 4

OpenStudy (debbieg):

If you distribute the -2 through on the left side, what do you get?

OpenStudy (anonymous):

x+a=-2x-4?

OpenStudy (debbieg):

Do you know what I mean when I say "distribute the -2 through"? The number outside the parentheses is multiplied THROUGH the parentheses.... e.g.... a(b + c) = ab + ac 5(y + 1) = 5y + 5(1) = 5y + 5

OpenStudy (debbieg):

Now, do that for -2(x + a) And what do you get?

OpenStudy (anonymous):

-4?

OpenStudy (debbieg):

Where did the variables go?? Look at my example above.

OpenStudy (debbieg):

a(b + c) = ab + ac 5(y + 1) = 5y + 5(1) = 5y + 5